OpenStudy (anonymous):

If cosx=4/5,cscx<0, then sin2x= cos2x= tan2x=

OpenStudy (anonymous):

if cos x= 4/5 sinx = 3/5

OpenStudy (anonymous):

sin2x = 2 sinx cosx 2 *4/5 * 3/5 = 24/25

OpenStudy (anonymous):

cos2x = 1- 2cos^x

OpenStudy (anonymous):

tan2x = sin2x /coss2x

OpenStudy (anonymous):

easy to calculate the rest